2/3+3/4=? 1/4+5/6=? 3/4+1/5=? 2/3+5/9=? 1/9+5/6=? 2/3+4/5=? 1/3+7/8=? 3/5+9/10=? 1/6+7/10=? 5/8+11/12=? 5/6+ 1/2=? 7/10+7/15=? 3/4+1/10=? 1/2+5/7=? I'm looking for the ansewers to these please?
Answer by edjones(7311)

We don't do your homework.
We show YOU how to do your homework.
.
2/3+3/4 What is the smallest number both 3 and 4 go into? 12
So the bottom number will be 12.
To find the top number for 2/3 divide the 3 into 12 and multiply the result times 2. answer: 8
With 3/4 divide 4 into 12 and multiply the result times 3 for an answer of 9.
8/12 + 9/12 = 17/12
.
We will do one more.
2/3+5/9 Bottom number (denominator) is 27
18/27 + 15/27 = 33/27= 11/9
.
Ed